OpenStudy (anonymous):

someone help? (:

OpenStudy (anonymous):

\[\int\limits_{-2}^{0} (\frac{ 1 }{ 2 }t^4+\frac{ 1 }{ 4 }t^3-t)dt\]

OpenStudy (nubeer):

Hint. \[\int\limits_{?}^{?} t ^{4} =\frac{ t ^{4+1} }{ 4+1}\]

OpenStudy (anonymous):

I honestly don't even know how to solve or begin to solve for it.. can you help me step by step? I have 7 more problems like this and I'd like to do them all myself.. But I need help solving and the steps to do so.

OpenStudy (anonymous):

are you familiar with the general formula of integration?

OpenStudy (anonymous):

I'm familiar with... \[\int\limits_{a}^{b} f(x)dx=F(x)|_{a}^{b} = F(b)-F(a)\]

OpenStudy (nubeer):

ya this is done when we apply the limits but do you know how we do integration?

OpenStudy (anonymous):

I don't.

OpenStudy (anonymous):

the general formula for integration of \(\int x^n dx = \frac{x^n+1}{n+1} +C\)

OpenStudy (nubeer):